OpenStudy (anonymous):

a box is pushed 40 m by a mover. the amount of work done was 2,240j. how much force was exerted on the box a box is pushed 40 m by a mover. the amount of work done was 2,240j. how much force was exerted on the box @Physics

OpenStudy (anonymous):

you just Isolate F of the formula W = F * D --> F = W/Dso F = 56N

OpenStudy (anonymous):

w=f.d => 2240 = f.40 => f=56N